OpenStudy (anonymous):

A body travels 10 meters during the first 5 seconds of its travel, and it travels a total of 30 meters over the first 10 seconds of its travel. What is its average speed during the time from t = 5 seconds to t = 10 seconds?

OpenStudy (anonymous):

2 meters/second 3 meters/second 4 meters/second 5 meters/second

OpenStudy (anonymous):

\[V =\frac{ \Delta x }{ \Delta t } = \frac{ x _{f}-x _{i} }{ t _{f}-t _{i} }=\frac{ 30-10 }{ 10-5 }= 4 m/s\]

OpenStudy (anonymous):

thank you sir